A) Banana, Inc. has a book value per share of $8.70, earnings per share of $1.68, and a price-earnings ratio of 22.2. What is the market-to-book ratio? B) Melon, Inc. has inventory of $2,200, current liabilities of $5,600, cash of $1,900, and accounts receivable of $3,900. What is the cash ratio?
Answer #1

Market to Book Ratio = Market price of stock / Book value of stock

Market price of stock = Price earning ratio * earnings per share

= 22.20 * 1.68 = 37.296

Market to book ratio = 37.296 / 8.70

= 4.29

Part B

Cash ratio = (Cash + cash equivalents) / current liabilities

= 1900 / 5600 = 0.34